OpenStudy (anonymous):

Solve 3(x – 5) – 15 = 48? totally forgot how to do these.. is it (3x-15)-15=48 then solve from there or what?

OpenStudy (anonymous):

Yes, then you solve from there, by adding 15 to both sides. And then adding 15 again and then dividing everything by 3.

OpenStudy (anonymous):

Do the opposite of PEMDAS. 3(x - 5) - 15 = 48 3(x - 5) = 63 x - 5 = 21 x = 26 If you do it the way you started: 3(x - 5) - 15 = 48 3x - 15 - 15 = 48 3x - 30 = 48 3x = 78 x = 26

OpenStudy (anonymous):

ah okay was doing it here beat me to it thank you